# Probing displaced (dark)photons from low reheating freeze-in at the LHC

### v1

@@ 4.1.3 Bound from a fit to combined measurements
The CMS CMS:2018uag and ATLAS ATLAS:2019nkf ; ATLAS:2022vkf collaborations have performed a likelihood fit of a combination of their data on Higgs production and decay, in order to constrain the Higgs couplings and its branching ratios into invisible and undetected particles.
In our case, since $\phi$ does not mix with the Higgs, the only deviation with respect to the SM Higgs properties comes from exotic decays. In particular, since displaced photon searches were not taken into account in the fit, our scenario is subject to the bound on the Higgs branching ratio to undetected particles $B_{\rm und}<0.12$ ATLAS:2022vkf . This bound is also reinterpreted, for each parameter space point $(m_{\gamma^{\prime}},m_{\phi})$ using eq. (6), as a bound on $\lambda_{HS}$. Note that it is independent of the value of the $g_{D}$ coupling.
@@ 4.2 Collider results and consequences on dark matter
[…]
@@
For both benchmark $\phi$ masses, we find that the displaced-photon search has the highest sensitivity for $c\,\tau_{\phi}\approx 2$ m. For this decay length, branching ratios down to about $0.5\%$ can be probed for $m_{\phi}=60$ GeV. This sensitivity is lower for $m_{\phi}=30$ GeV, being able to constrain branching ratios around $7\%$. The increased sensitivity for larger masses was explained in Duarte:2023tdw . On the one hand, larger masses lead to slower moving LLPs, such that the displaced photon arrives at the ECal with a larger delay. On the other hand, with a smaller boost the momenta of the LLP decay products are much less collimated, so the displaced photon points less towards the primary vertex, leading to a larger $|\Delta z_{\gamma}|$. Thus, LLPs with larger masses have a greater probability that their displaced photons will be assigned to the signal region of the search.
@@
We present our final result on the left panel of figure 9, as contour lines giving the maximum allowed $\lambda_{HS}$ in the $g_{D}-m_{\phi}$ space, for $m_{\gamma^{\prime}}=1$ GeV. The contours in the middle right zone of the plot, with colors from red to light blue, constrain $\lambda_{HS}$ to values below 22225.$5.5\times 10^{-3}$, and come from the search for displaced photons. On the other hand, the vertical lines on the upper right, with colors in blue, gray and dark-cyan, stem from the Higgs to undetected bound and, as we mentioned earlier, are independent of the value of $g_{D}$. The rest of the contours, on the left and lower parts of the plot, are governed by the Higgs to invisible bound, following eq. (28). These contours are consistent with our results in figure 8, and demonstrate the synergies of the three different analyses when constraining the parameter space.
@@ 6 Acknowledgments
We thank Sasha Pukhov for his constant help with micrOMEGAs. J.J.P., W.R. and D.Z. acknowledge financial support from the Dirección de Fomento de la Investigación at PUCP, though grant DFI-PUCP-PI1144. BDS was funded by ANID FONDECYT postdoctorado 2022 No. 3220566 and ANID FONDECYT Iniciación No. 1125181. P.A. acknowledges support from FONDECYT project Nº 1251613. L.D. acknowledges support from PEDECIBA (Uruguay).
@@ B.1 LHC signal simulation and improved non-pointing variable
In our signal simulation, we consider that $\phi$ is pair-produced promptly in proton-proton collisions as decay products of on-shell Higgs bosons $h$. Then, each long-lived scalar decays to one dark photon $\gamma^{\prime}$ and a photon $\gamma$ that can be detected at the ECal. These photons, coming from an LLP, can lead to a non-pointing photon signal. We considered the triggering leptons coming from $p\,p\to W^{\pm}h$, $p\,p\to Z\,h$ and $p\,p\to t\,\bar{t}\,h$ processes. The model was built in FeynRules 2.3.43 Christensen:2008py ; Alloul:2013bka , and events were generated with MadGraph5_aMC@NLO 2.9.7 Alwall:2014hca , which uses LHAPDF6 Buckley:2014ana . PYTHIA 8.244 Sjostrand:2006za performs the $\phi$ scalar decay, parton showering and hadronization.
The interested reader can find details of how $t_{\gamma}$ was modeled in Delgado:2022fea ; Duarte:2023tdw . However, for this work we carry out a refined calculation of $|\Delta z_{\gamma}|$. The non-pointing variable used in ATLAS:2022vhr is defined as Nikiforou:2014cka ; Mahon:2021bai :
@@
where $z_{i},\,R_{i}$ are the coordinates of the position of the centroid of the electromagnetic shower associated to each photon on the first two layers of the detector, as illustrated in figure 7. The dependence of the radial depths $R_{i}$ on the reconstructed pseudorapidity are shown on page 48 (figure 5) in ATLAS:2009zsq .
For this work we improve the calculation of $|\Delta z_{\gamma}|$ by using the information on $R_{i}$ found in ATLAS:2009zsq . First, we use the truth-level vectors $\vec{r}_{\phi}$ and $\vec{p}_{\gamma}$, for the $\phi$ decay position and momentum of its daughter photon, respectively, to describe the photon’s trajectory inside the detector:
@@
where $\alpha$ is an evolution parameter. Using this equation, one can calculate the $z_{1}$ and $z_{2}$ coordinates where $\vec{L}$ intersects $r=R_{1}$ and $r=R_{2}$ surfaces, for specific $R_{i}$, regardless of the azimuthal angle $\varphi$. In order to determine which value of $R_{i}$ we should use, we first find the pseudorapidity $\eta_{i}$ of the photon as it enters each layer, at $r=1500$ and $1590$ mm, respectively (see figure 5.4 of TheATLASCollaboration_2008 ). Notice that this pseudorapidity refers to a region of the detector, as measured from the interaction point. So, for each value of $\eta_{i}$, we find the appropriate $R_{i}$ from the corresponding curves of figure 5 of ATLAS:2009zsq . Using then eqs. (33) and (34), we obtain:
@@
We have compared this new method with the old one used in Delgado:2022fea ; Duarte:2023tdw , which performs a 3D reconstruction of $|\Delta z_{\gamma}|$ instead of the 2D used by ATLAS. The difference is given by the last term in eq. (35). We found that large part of the effect are washed out by the detector resolution. However, it is important to clarify that we did find events where the old $|\Delta z_{\gamma}|$ differed significantly with respect to the new one. Fortunately, these were not common, and did not have significant impact on the final results in Delgado:2022fea ; Duarte:2023tdw .
@@ B.2 Event reconstruction and improved muon resolution
The event reconstruction is simulated by Delphes 3.5.0 deFavereau:2013fsa , which depends on FastJet 3.4.0 Cacciari:2011ma . The detector simulation includes most modules from the ATLAS card packaged within Delphes, with modifications as described in Duarte:2023tdw . In particular, we adapted photon/electron track isolation, photon/electron calorimeter isolation, and photon/electron ID efficiencies, according to ATLAS:2022vhr .
In this work, we have improved the muon isolation implementation compared to Duarte:2023tdw , in order to apply an overlap removal algorithm closer to that in ATLAS:2022vhr . Our procedure for muon isolation is now based on information from the PflowLoose working point of section 5.3 in ATLAS:2020auj , as found in Chapter 6.2.2 in Mahon:2021bai . For definiteness, we consider tracks with transverse momentum larger that 0.5 GeV, and all energy deposits from neutral particles, within a $\Delta R=0.2$ fixed cone around each muon. The scalar sum of these tracks is labelled $p_{\rm trk}$, while that of energy deposits is $E_{T}^{\rm neut}$. From this, a muon with transverse momentum $p_{T}^{\mu}$ is considered as isolated if $(p_{\rm trk}+0.4E_{T}^{\rm neut})<0.16\,p_{T}^{\mu}$. Our procedure is implemented by appropriately modifying the corresponding Delphes isolation subroutine, using the tracks and eflowNeutralHadrons variables to calculate $p_{\rm trk}$ and $E_{T}^{\rm neut}$, respectively. Since muon isolation has an efficiency above $95\%$, we have introduced this new procedure without modifying the total identification efficiency used in Duarte:2023tdw .
After object reconstruction it is necessary to implement the overlap removal algorithm. According to ATLAS:2022vhr , if an electron or jet is found within $\Delta R\leq 0.4$ of a photon, they are removed from the event. Similarly, if a jet is found within $\Delta R\leq 0.2$ from an electron, it is also removed from the event. However, if an electron is found within $\Delta R\leq 0.4$ from any surviving jet, then the electron is removed. Finally, any muon with $\Delta R\leq 0.4$ from photons or surviving jets were removed.
The last step of this algorithm, i.e. favouring jets over muons, provided problematic, as Delphes initially classifies all muons as jets, and then performs its own overlap removal on the UniqueObjectFinder module, which favours muons over jets. In Duarte:2023tdw , an independent overlap removal was performed, but faced the problem of having all muons classified as jets by Delphes. To overcome this, we modified the Isolation module in Delphes to return both isolated and non-isolated muons, allowing us to consider only the non-isolated ones in the jet algorithm.
@@ B.3 Statistical analysis
We calculate our 95% exclusion limits using the CLs method, as done in Duarte:2023tdw . The upper limit to the number of signal events $s^{\rm up}$ is $6.8$ ($3.8$) for the single (multi) photon channels. The search in ATLAS:2022vhr also presents a “combination” of both channels, with $s^{\rm up}=6.8$. However, this is essentially a combination of the information from the single and multi-photon channels into a single category, which reflects the fact that both channels are maximally correlated due to the procedure for assigning the events to one or the other category.
@@
where $\mathcal{L}$ is the integrated luminosity, $\sigma_{Xh}$ are the cross-sections for each Higgs production process, $K_{X}$ are the associated K-factors LHCHiggsCrossSectionWorkingGroup:2016ypw , $N_{X}^{\rm gen}$ are the number of generated events for each process, and $N_{X}^{\rm cuts}$ are the corresponding number of events surviving the cuts. This bound on ${\rm BR}(h\to\phi\phi)$ can also be translated into a limit on $\lambda_{HS}$, using eq. (6).

### v2

@@ 4.1.3 Bound from a fit to combined measurements
The CMS Sirunyan and others (2019a) and ATLAS Aad and others (2020, 2022) collaborations have performed a likelihood fit of a combination of their data on Higgs production and decay, in order to constrain the Higgs couplings and its branching ratios into invisible and undetected particles.
In our case, since $\phi$ does not mix with the Higgs, the only deviation with respect to the SM Higgs properties comes from exotic decays. In particular, since displaced photon searches were not taken into account in the fit, our scenario is subject to the bound on the Higgs branching ratio to undetected particles $B_{\rm und}<0.12$ Aad and others (2022). This bound is also reinterpreted, for each parameter space point $(m_{\gamma^{\prime}},m_{\phi})$ using eq. (6), as a bound on $\lambda_{HS}$. Note that it is independent of the value of the $g_{D}$ coupling.
@@ 4.2 Collider results and consequences on dark matter
Figure 8: Left: Constraints of ${\rm BR}(h\to\phi\phi)$ as a function of $c\,\tau_{\phi}$. Right: Constraints of $\lambda_{HS}$ as a function of $g_{D}$. In both panels, we present bounds coming from displaced photons (solid), Higgs invisible decays (dashed) and Higgs decays to undetected particles (shaded regions). Benchmarks points with $m_{\phi}=30,\,60$ GeV are shown in blue and orange, respectively, both having $m_{\gamma^{\prime}}=1$ GeV.
@@
For both benchmark $\phi$ masses, we find that the displaced-photon search has the highest sensitivity for $c\,\tau_{\phi}\approx 2$ m. For this decay length, branching ratios down to about $0.5\%$ can be probed for $m_{\phi}=60$ GeV. This sensitivity is lower for $m_{\phi}=30$ GeV, being able to constrain branching ratios around $7\%$. The increased sensitivity for larger masses was explained in Duarte et al. (2024). On the one hand, larger masses lead to slower moving LLPs, such that the displaced photon arrives at the ECal with a larger delay. On the other hand, with a smaller boost the momenta of the LLP decay products are much less collimated, so the displaced photon points less towards the primary vertex, leading to a larger $|\Delta z_{\gamma}|$. Thus, LLPs with larger masses have a greater probability that their displaced photons will be assigned to the signal region of the search.
@@
[…]
@@ 6 Acknowledgments
We thank Sasha Pukhov for his constant help with micrOMEGAs. J.J.P., W.R. and D.Z. acknowledge financial support from the Dirección de Fomento de la Investigación at PUCP, though grant DFI-PUCP-PI1144. B.D.S. was funded by ANID FONDECYT postdoctorado 2022 No. 3220566 and ANID FONDECYT Iniciación No. 1125181. B.D.S. also acknowledges support from ANID — Millennium Science Initiative Program ICN2019_044 . P.A. acknowledges support from FONDECYT project Nº 1251613. L.D. acknowledges support from PEDECIBA (Uruguay).
@@ B.1 LHC signal simulation and improved non-pointing variable
In our signal simulation, we consider that $\phi$ is pair-produced promptly in proton-proton collisions as decay products of on-shell Higgs bosons $h$. Then, each long-lived scalar decays to one dark photon $\gamma^{\prime}$ and a photon $\gamma$ that can be detected at the ECal. These photons, coming from an LLP, can lead to a non-pointing photon signal. We considered the triggering leptons coming from $p\to W^{\pm}h, pp$$p\to Z\,h$ and $p\,p\to t\,\bar{t}\,h$ processes. The model was built in FeynRules 2.3.43 Christensen and Duhr (2009); Alloul et al. (2014), and events were generated with MadGraph5_aMC@NLO 2.9.7 Alwall et al. (2014), which uses LHAPDF6 Buckley et al. (2015). PYTHIA 8.244 Sjostrand et al. (2006) performs the $\phi$ scalar decay, parton showering and hadronization.
The interested reader can find details of how $t_{\gamma}$ was modeled in Delgado et al. (2022); Duarte et al. (2024). However, for this work we carry out a refined calculation of $|\Delta z_{\gamma}|$. The non-pointing variable used in Aad and others (2023b) is defined as Nikiforou (2014); Mahon (2021):
@@
where $z_{i},\,R_{i}$ are the coordinates of the position of the centroid of the electromagnetic shower associated to each photon on the first two layers of the detector, as illustrated in figure 7. The dependence of the radial depths $R_{i}$ on the reconstructed pseudorapidity are shown on page 48 (figure 5) in Aad and others (2009).
For this work we improve the calculation of $|\Delta z_{\gamma}|$ by using the information on $R_{i}$ found in Aad and others (2009). First, we use the truth-level vectors $\vec{r}_{\phi}$ and $\vec{p}_{\gamma}$, for the $\phi$ decay position and momentum of its daughter photon, respectively, to describe the photon’s trajectory inside the detector:
@@
where $\alpha$ is an evolution parameter. Using this equation, one can calculate the $z_{1}$ and $z_{2}$ coordinates where $\vec{L}$ intersects $r=R_{1}$ and $r=R_{2}$ surfaces, for specific $R_{i}$, regardless of the azimuthal angle $\varphi$. In order to determine which value of $R_{i}$ we should use, we first find the pseudorapidity $\eta_{i}$ of the photon as it enters each layer, at $r=1500$ and $1590$ mm, respectively (see figure 5.4 of Aad and others (2008)). Notice that this pseudorapidity refers to a region of the detector, as measured from the interaction point. So, for each value of $\eta_{i}$, we find the appropriate $R_{i}$ from the corresponding curves of figure 5 of Aad and others (2009). Using then eqs. (33) and (34), we obtain:
@@
We have compared this new method with the old one used in Delgado et al. (2022); Duarte et al. (2024), which performs a 3D reconstruction of $|\Delta z_{\gamma}|$ instead of the 2D used by ATLAS. The difference is given by the last term in eq. (35). We found that large part of the effect are washed out by the detector resolution. However, it is important to clarify that we did find events where the old $|\Delta z_{\gamma}|$ differed significantly with respect to the new one. Fortunately, these were not common, and did not have significant impact on the final results in Delgado et al. (2022); Duarte et al. (2024).
@@ B.2 Event reconstruction and improved muon resolution
The event reconstruction is simulated by Delphes 3.5.0 de Favereau et al. (2014), which depends on FastJet 3.4.0 Cacciari et al. (2012). The detector simulation includes most modules from the ATLAS card packaged within Delphes, with modifications as described in Duarte et al. (2024). In particular, we adapted photon/electron track isolation, photon/electron calorimeter isolation, and photon/electron ID efficiencies, according to Aad and others (2023b).
In this work, we have improved the muon isolation implementation compared to Duarte et al. (2024), in order to apply an overlap removal algorithm closer to that in Aad and others (2023b). Our procedure for muon isolation is now based on information from the PflowLoose working point of section 5.3 in Aad and others (2021), as found in Chapter 6.2.2 in Mahon (2021). For definiteness, we consider tracks with transverse momentum larger that 0.5 GeV, and all energy deposits from neutral particles, within a $\Delta R=0.2$ fixed cone around each muon. The scalar sum of these tracks is labelled $p_{\rm trk}$, while that of energy deposits is $E_{T}^{\rm neut}$. From this, a muon with transverse momentum $p_{T}^{\mu}$ is considered as isolated if $(p_{\rm trk}+0.4E_{T}^{\rm neut})<0.16\,p_{T}^{\mu}$. Our procedure is implemented by appropriately modifying the corresponding Delphes isolation subroutine, using the tracks and eflowNeutralHadrons variables to calculate $p_{\rm trk}$ and $E_{T}^{\rm neut}$, respectively. Since muon isolation has an efficiency above $95\%$, we have introduced this new procedure without modifying the total identification efficiency used in Duarte et al. (2024).
After object reconstruction it is necessary to implement the overlap removal algorithm. According to Aad and others (2023b), if an electron or jet is found within $\Delta R\leq 0.4$ of a photon, they are removed from the event. Similarly, if a jet is found within $\Delta R\leq 0.2$ from an electron, it is also removed from the event. However, if an electron is found within $\Delta R\leq 0.4$ from any surviving jet, then the electron is removed. Finally, any muon with $\Delta R\leq 0.4$ from photons or surviving jets were removed.
The last step of this algorithm, i.e. favouring jets over muons, provided problematic, as Delphes initially classifies all muons as jets, and then performs its own overlap removal on the UniqueObjectFinder module, which favours muons over jets. In Duarte et al. (2024), an independent overlap removal was performed, but faced the problem of having all muons classified as jets by Delphes. To overcome this, we modified the Isolation module in Delphes to return both isolated and non-isolated muons, allowing us to consider only the non-isolated ones in the jet algorithm.
@@ B.3 Statistical analysis
The ATLAS search estimated the dominant background to be due to pp collisions. The measured number of events was consistent with the background expectation, implying constraints on the production of signal events. We calculate our 95% exclusion limits using the CLs method, as done in Duarte et al. (2024). The upper limit to the number of signal events $s^{\rm up}$ is $6.8$ ($3.8$) for the single (multi) photon channels. The search in Aad and others (2023b) also presents a “combination” of both channels, with $s^{\rm up}=6.8$. However, this is essentially a combination of the information from the single and multi-photon channels into a single category, which reflects the fact that both channels are maximally correlated due to the procedure for assigning the events to one or the other category.
@@
where $\mathcal{L}$ is the integrated luminosity, $\sigma_{Xh}$ are the cross-sections for each Higgs production process, $K_{X}$ are the associated K-factors de Florian and others (2016), $N_{X}^{\rm gen}$ are the number of generated events for each process, and $N_{X}^{\rm cuts}$ are the corresponding number of events surviving the cuts. This bound on ${\rm BR}(h\to\phi\phi)$ can also be translated into a limit on $\lambda_{HS}$, using eq. (6).